1. Find a homogeneous linear differential equation with constant coefficients whose general solution is given.

y = c1e−x cos x + c2e−x sin x

A. y'' + 2y' + 2y = 0
B. y'' − 2y' + 2y = 0
C. y''' + 2y'' + 2y' = 0
D. y''' − 2y'' + 2y' = 0
E. y'' + 1 = 0

2. Find the general solution of the given second-order differential equation.
y'' + 36y = 0

y(x) = ____
1
Expert's answer
2020-07-05T17:38:57-0400

1.

Solution:

y = c1e-xcos x + c2-x sin x,

k1=-1-i, k2=-1+i,

(k-k1)(k-k2)=(k+1+i)(k+1-i)=k2+2k+2,

y''+2y'+2y=0

Answer:

A.y''+2y'+2y=0.


2.

Solution:

y''+36y=0,

k2+36=0,

k2=-36,

k1=-6i, k2=6i,

y(x)=c1cos(6x)+c2sin(6x).

Answer:

y(x)=c1cos(6x)+c2sin(6x).